S Company's net income was earned evenly throughout the year. Both companies declared and paid their dividends on December 31, 20X8. P uses the fully adjusted equity method in accounting for its investment in S.

P Corporation acquired 70 percent ownership of Company on January 1, 20X6, at underlying book value. At that date, the fair vPlease help me prepare the elimination entries needed to complete full consolidation worksheet for 20X8
Answer #1
Date Particulars Debit ($) Credit ($)
Dividend income      14,000
Non controlling interest of S co.        6,000
Dividend declared           20,000
(To record the dividend consolidation entry)
Date Particulars Debit ($) Credit ($)
Common stock      60,000
Retained earnings      90,000
Investment in S Co. Stock         105,000
Non controlling interest of           45,000
(To record investment consolidation entry)
Date Particulars Debit ($) Credit ($)
Bonds payable    100,000
Loss on bonds retirement      18,200
Discount on bond payable             4,200
Investment in S co.         114,000
(To record the elimination of inter company bond holding)

Explanation

Calculations:

Date Particulars Debit ($) Credit ($)
Dividend income =20000*70%
Non controlling interest of S co. =20000*30%
Dividend declared 20000
(To record the dividend consolidation entry)
Date Particulars Debit ($) Credit ($)
Common stock =50000+10000
Retained earnings 90000
Investment in S Co. Stock =150000*70%
Non controlling interest of =150000*30%
(To record investment consolidation entry)
Date Particulars Debit ($) Credit ($)
Bonds payable 100000
Loss on bonds retirement =+G22+G23-F20
Discount on bond payable =+(114000-100000)*30%
Investment in S co. 114000
(To record the elimination of inter company bond holding)